# Improving Sequential Market Coordination via Value-oriented Renewable Energy Forecasting

### v2

@@ Abstract
Keywords: Energy forecasting, Loss function, Forecast value, Market clearing, Decision-focused learning
@@ I Introduction
A significant challenge with sequential deterministic market clearing arises from its limited ability to coordinate the two markets under large RES uncertainty. Currently, the DA market is cleared without accounting for its impact on RT market clearing. As a result, DA and RT markets operate without coordination, leading to high overall operating costs [3]. Consequently, stochastic market clearing has been proposed, which informs the DA market of the operating cost in the RT market [4]. While stochastic market clearing can improve overall economic efficiency, it struggles to simultaneously maintain key market properties, such as revenue adequacy and cost recovery [5]. To address this, efforts have been made to preserve these desirable properties within stochastic market clearing. For instance, [6] guarantees cost recovery and revenue adequacy both per scenario and in expectation, though this comes at the cost of reduced market efficiency.
Previous studies have shown that combining DA deterministic market clearing with RES improved entering quantities (RIEQ) enhances coordination between DA and RT markets, achieving performance comparable to stochastic clearing [7, 8]. The key idea is to tactically determine RIEQ—used as input parameters in DA clearing—by considering their impact on future RT market operations. For instance, studies [7, 8] optimize RIEQ by solving a bilevel program on a case-by-case basis during the operational phase. Similarly, [9] applies these methods to predict reserve requirements, facilitating coordination of reserve and energy markets. While these approaches hold promise for improving market coordination, they introduce computational challenges in determining RIEQ.
Building on this foundation, the goal of this work is not to replace the existing RES entering quantities determined by RES producer offerings in the DA market. Instead, we aim to explore an important question: What if the RES entering quantities in the DA market were improved as RIEQ by a “properly” trained forecasting model, specifically optimized to minimize overall system operating costs? If so, could the sequential markets be better coordinated? Here, the overall system operating cost in both the DA and RT markets is used as the quantitative metric to evaluate market coordination.
@@
For the first thread of research on integrated optimization, forecasting model parameters are optimized concurrently with decision variables [11, 13]. In the same vein, [12] introduces a bilevel program where DA market clearing forms the lower level with the RIEQ prediction as a parameter, while the upper level optimizes both model parameters and RT decisions. This method requires a linear forecasting model to make the program solvable by commercial optimization solvers, which may limit the performance. The second thread of research, on differentiable programming, enables the use of more advanced forecasting models, such as neural networks, by deriving the gradient of decision solution w.r.t. the forecast [14, 15]. However, obtaining such a gradient involves repeatedly solving the inverse of a large-size matrix resulting from the Karush–Kuhn–Tucker conditions, making it computationally expensive. The third approach, and the focus of our work, centers on designing value-oriented loss functions. The existing ones, such as pinball loss [16] or Smart “Predict, then Optimize” (SPO) loss [17], are primarily designed for single-stage stochastic programs (e.g., the one without considering flexible unit redispatch in the RT market). Thus, how to develop a value-oriented loss function for sequential market clearing remains an open question.
@@
To address the challenge of training the RIEQ forecasting model, we propose a tailored loss function that uses the RIEQ forecast as the input and the overall DA and RT system operating costs as the output. The forecasting model parameter estimation is formulated as a bilevel program: the upper level optimizes forecasting model parameters, while the lower level solves the DA and RT market clearing problems using the RIEQ forecast as the input parameter. To link operating cost with the RIEQ forecast, we leverage the lower-level dual problems and replace the upper-level objective with dual objectives. The reformulated upper-level objective is then transformed into an analytical function of the forecast, serving as the loss function for training. This requires the derivation of functions that connect the forecast with primal and dual solutions. Specifically, the dual solutions are linked to forecasts through the dual problems of DA and RT market clearing, while the primal solutions are derived from the active constraints of their respective primal problems. Substituting these functions into the upper-level objective yields a value-oriented loss function for training. Our main contributions are,
@@
1) From a market perspective, the proposed approach maintains deterministic DA market clearing while improving sequential market coordination with RIEQ forecasts.
2) From a theoretical perspective, we derive a value-oriented loss function that aligns the forecasting model’s training objective with operational value, minimizing overall system operating costs in both DA and RT markets.
3) From a practitioner’s perspective, the analytical loss function allows the analytical derivation of the gradient. Leveraging the structure of the gradient, we propose a computationally efficient training approach.
Although prior works [18, 19] design forecasting products using value-oriented loss functions, these approaches are not directly applicable to electricity market settings. First, prior works cannot address the participation of entities other than RES in both DA and RT markets, which is typical in electricity markets. As a result, the network and the redispatch of flexible resources cannot be handled. Second, the generalized market setting introduces new technical challenges in obtaining the value-oriented loss function. The reformulated upper-level objective, derived from dual problems, depends on both the forecast and the primal and dual solutions from the lower level. This requires us to identify the functions between those solutions and the forecast.
@@ II Preliminaries
Figure 1: An illustration of forecasting and the sequential market clearing where $\hat{\bm{y}}_{d,\tau}$ and $\bm{y}_{d,\tau}$ are forecasts and realization, and $\bm{w}_{d,\tau}^{*}$ is the RES schedule.
@@ II-A Sequential Market Clearing
We consider the sequential clearing of DA and RT markets [1], as illustrated in Fig. 1. The DA market is cleared at time $t$ on day $d-1$, with an advance of $k$ hours in time to the next day $d$, and covers energy transactions on day $d$, typically on an hourly basis. RES production is uncertain during the DA market, where the RES forecast acts as the RIEQ. Due to inevitable forecasting errors, the energy imbalance caused by forecasting errors needs to be settled in the RT market. Concretely, in the DA market, the operator determines the schedules of generators and RES to satisfy inelastic demand. The generation and RES schedule for each time-slot $\tau,\forall\tau=1,...,24$ on the next day $d$ are denoted as $\bm{p}_{d,\tau}$ and $\bm{w}_{d,\tau}$, respectively. The DA market clearing is,
@@
s.t. $\bm{1}^{\top}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau})=\bm{1}^{\top}\bm{l}_{d,\tau},\forall\tau=1,...,T$ (1b)
$-\overline{\bm{f}}\leq\bm{H}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau}-\bm{l}_{d,\tau})\leq\overline{\bm{f}},\forall\tau=1,...,T$ (1c)
$\underline{\bm{p}}\leq\bm{p}_{d,\tau}\leq\overline{\bm{p}},\forall\tau=1,...,T$ (1d)
$\underline{\bm{r}}\leq\bm{p}_{d,\tau}-\bm{p}_{d,\tau-1}\leq\overline{\bm{r}},\forall\tau=2,...,T$ (1e)
@@
In line with European practices [12, 21], we do not incorporate binary decisions regarding unit commitment (UC) in (1). However, we note that UC is a requisite consideration in the U.S. markets. To analyze the market behavior, the relaxed UC problem, where the binary commitment decisions are substituted with continuous ones, is widely used [3, 8]. The relaxed UC is a linear program. A common practice is to train the forecasting model using the relaxed UC formulation [8]. During the operational phase, the generated forecasts are then used as inputs to the UC problem.
Since the RES production is uncertain in DA, the DA schedules are to be adjusted at each time-slot $\tau,\forall\tau=1,...,T$ in RT on day $d$, after the RES realization $\bm{y}_{d,\tau}$ is observed. The RT market deals with the imbalance $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ caused by RES, with a minimized imbalance cost. Additionally, the RT market clearing at time-slot $\tau,\forall\tau=2,...,T$ is influenced not only by the DA clearing outcomes but also by the extent of power adjustments made in the preceding time-slot. This is due to the ramping constraints that interconnect adjacent time slots. Because the market clearing is conducted separately for each day, the RT market clearing at time-slot $\tau=1$ on day $d$ remains unaffected by any adjustments made at time-slot $\tau=T$ on the previous day $d-1$. In the following, we firstly give the mathematical formulation of the RT clearing at $\tau=1$,
@@ II-B Mathematical Reformulation
$\underline{\bm{p}}-\bm{p}_{d,\tau}^{*}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{p}}-\bm{p}_{d,\tau}^{*}$ (5a)
$\bm{p}_{d,\tau}^{*}-\overline{\bm{p}}\leq\bm{p}_{d,\tau}^{-}\leq\bm{p}_{d,\tau}^{*}-\underline{\bm{p}}$ (5b)
Since $\underline{\bm{p}}\leq\bm{p}_{d,\tau}^{*}\leq\overline{\bm{p}}$, the left side of (5) is less than 0. Considering the power adjustment $\bm{p}_{d,\tau}^{+},\bm{p}_{d,\tau}^{-}$ is larger than 0 as stated in (2d) and (2e), (5) can be further simplified as,
@@
III Methodology: Parameter Estimation
Figure 2: An illustration of the bilevel program.
In this section, we show how to identify the RIEQ, whose value corresponds to $\hat{\bm{y}}_{d,\tau}$, an input parameter in (1), via training forecasting models. At the training phase, a bilevel program [18] is formulated for estimating the forecasting model parameter. Let $g(\ \cdot\ ;\Theta)$ denote the forecasting model with the parameter $\Theta$, and $\bm{s}_{d,\tau}$ denote the context. The RIEQ forecast for the time $\tau$ on day $d$ is,
@@ III Methodology: Parameter Estimation
where $\hat{\bm{y}}_{d,\tau}$ denotes the forecast at the training phase, with $\Theta$ given by any value. Data in the training set $\{\{\bm{s}_{d,\tau},y_{d,\tau}\}_{\tau=1}^{T}\}_{d=1}^{D}$ is available, which is consisted of historical context and RES realization in $D$ days. An illustration of the bilevel program is shown in Fig. 2. The upper level determines the model parameter $\Theta$, while the lower level involves the DA and RT market clearings. The bilevel program is mathematically formulated as,
@@
(12)$,\forall d=1,...,D$ (15i)
Since RT clearing requires the primal solutions of DA clearing and the previous RT clearing as input parameters, we also include the primal problems in the lower level. The forecast $\hat{\bm{y}}_{d}$ affects the upper-level objective (15a) via its impact on the DA and RT dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*}$, and their primal solutions $\bm{x}^{*}_{d,\tau},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$. If we can obtain the function between them and the forecast $\hat{\bm{y}}_{d}$ directly, the upper-level objective can be rewritten as a function regarding the forecast $\hat{\bm{y}}_{d}$, and can be used as the loss function for training. In the next section, we will show how to achieve this.
At the operation phase, the forecast $\bm{\hat{y}}_{d,\tau}$ under the context $\bm{s}_{d,\tau}$ can be obtained by the trained model. Subsequently, utilizing this forecast, the operator proceeds to solve the DA clearing in (1), followed by the RT clearing in (2) and (3).

### v3

@@ Abstract
Keywords: Energy forecasting, Surrogate Loss, Forecast value, Market clearing, Decision-focused learning
@@ I Introduction
A significant challenge with sequential deterministic market clearing arises from its limited ability to coordinate the two markets under large RES uncertainty. Specifically, DA clearing decisions are made under RES uncertainty, influencing RT market clearing, which subsequently manages power imbalances arising from discrepancies between RES schedules and realizations. Currently, the DA market is cleared without accounting for its impact on RT market clearing. As a result, DA and RT markets operate without coordination, leading to high overall operating costs [3]. Consequently, stochastic market clearing has been proposed, which informs the DA market of the operating cost in the RT market [4]. While stochastic market clearing can improve overall economic efficiency, it struggles to simultaneously maintain key market properties, such as revenue adequacy and cost recovery [5]. To address this, efforts have been made to preserve these desirable properties within stochastic market clearing. For instance, [6] guarantees cost recovery and revenue adequacy both per scenario and in expectation, though this comes at the cost of reduced market efficiency.
Previous studies have shown that combining DA deterministic market clearing with RES improved entering quantities (RIEQ) enhances coordination between DA and RT markets, achieving performance comparable to stochastic clearing [7, 8]. The key idea is to tactically determine RIEQ, which is made under uncertainty and used as input parameters in DA clearing, by considering the impact of DA clearing results on future RT market clearing. For instance, studies [7, 8] optimize RIEQ by solving a bilevel stochastic program on a case-by-case basis during the operational phase. Similarly, [9] applies these methods to predict reserve requirements, facilitating coordination of reserve and energy markets. While these approaches hold promise for improving market coordination, they introduce computational challenges in determining RIEQ.
Building on this foundation, we aim to explore an important question: What if the RES entering quantities in the DA market were improved as RIEQ by a “properly” trained forecasting model, specifically optimized to minimize overall system operating costs? If so, could the sequential markets be better coordinated? Here, the overall system operating cost in both the DA and RT markets is used as the quantitative metric to evaluate market coordination.
@@
For the first thread of research on integrated optimization, forecasting model parameters are optimized concurrently with decision variables [11, 13]. In the same vein, [12] introduces a bilevel program where DA market clearing forms the lower level with the RIEQ prediction as a parameter, while the upper level optimizes both model parameters and RT decisions. This method requires a linear forecasting model to make the program solvable by commercial optimization solvers, which may limit the performance. The second thread of research, on differentiable programming, enables the use of more advanced forecasting models, such as neural networks, by deriving the gradient of decision solution w.r.t. the forecast [14, 15]. However, obtaining such a gradient involves repeatedly solving the inverse of a large-size matrix resulting from the Karush–Kuhn–Tucker conditions, making it computationally expensive. In a general sense, the first two threads implicitly design value-oriented loss functions by aligning forecasting model training with decision-making value, without explicitly formulating a loss function that links forecasts to the decision-making objective. The third approach, and the focus of our work, centers on explicitly designing value-oriented loss functions. Compared to the implicit loss function, the explicit loss function has a clearer structure, which enhances the explainability of value-oriented forecasts [16]. The existing ones, such as pinball loss [17] or Smart “Predict, then Optimize” (SPO) loss [18], are primarily designed for single-stage stochastic programs (e.g., the one without considering flexible unit redispatch in the RT market). Thus, how to develop a value-oriented loss function for sequential market clearing remains an open question.
@@
To address the challenge of training the RIEQ forecasting model, we propose a tailored loss function that uses the RIEQ forecast as the input and the overall DA and RT system operating costs as the output. The forecasting model parameter estimation is formulated as a bilevel program: the upper level optimizes forecasting model parameters, while the lower level solves the DA and RT market clearing problems using the RIEQ forecast as the input parameter. To link operating cost with the RIEQ forecast, we leverage the lower-level dual problems and replace the upper-level objective with dual objectives. The reformulated upper-level objective is then transformed into an analytical function of the forecast, serving as the loss function for training. This requires the derivation of functions that connect the forecast with primal and dual solutions. Specifically, the dual solutions are linked to forecasts through the dual problems of DA and RT market clearing, while the functions for primal solutions are derived from the active constraints of their respective primal problems. Substituting these functions into the upper-level objective yields a value-oriented loss function for training. Our main contributions are twofold: from the market perspective and from the methodological perspective of obtaining RIEQ,
@@
1) From the market perspective, the proposed approach maintains deterministic DA market clearing while improving sequential market coordination with RIEQ forecasts. Additionally, we theoretically demonstrate that deterministic DA market clearing with RIEQ forecasts maintains key market properties, namely cost recovery and revenue adequacy.
2) From the methodological perspective of obtaining RIEQ, we propose determining RIEQ via a trained forecasting model at the operational phase. During the training phase, we analytically derive a value-oriented loss function that aligns the forecasting model’s training objective with the operational value, aiming to minimize overall system operating costs in both DA and RT markets. We reveal that the loss function is piecewise linear for market clearing problems modeled by linear programs, with each piece associated with active constraints. Leveraging insights from the loss function structure, we propose a computationally efficient training approach for gradient descent-based methods, which recalculates the gradient only when encountering new active constraints.
The above contributions distinguish this work from our prior study [16]. This work directly addresses the coordination challenge posed by RESs in electricity markets. Additionally, the method for deriving the loss function is specifically tailored to the market clearing problem and is more generalized compared to the approaches in [16]. Unlike the previous work [16], where the DA and RT operating problems are connected solely through the forecasts, in this work, the DA and RT market clearing are linked through the DA decisions, such as the schedules of traditional and renewable generators, which are implicitly influenced by the forecasts. This is specific to the market setting, as the schedules of flexible resources are determined in the DA market, while their flexibility enables them to make RT adjustments based on the DA schedules to correct RT power imbalances. The influence of the forecasts on the RT clearing is indirect, since RES forecasts are not explicitly included in the RT clearing process. The method for deriving the loss function in this work is specifically designed to account for this implicit impact, extending beyond the approach proposed in [16].
Also, it is worth mentioning that this work has not fully addressed all the challenges of implementing RIEQ forecasts in practical market clearing, such as how to align RIEQ with RES producer offering quantities. Our primary goal is to demonstrate that RIEQ can be determined via trained forecasting models at the operational phase, making it computationally convenient. Additionally, the RIEQ can enhance the coordination between day-ahead and real-time markets, thereby reducing overall operating costs while preserving key market properties, such as cost recovery and revenue adequacy.
@@
Table I: Nomenclature
Sets
$\mathcal{J}^{a}_{DA,d}$
Row index set of active constraints of day-ahead clearing
on day $d$
$\mathcal{J}^{a}_{RT,d,\tau}$
Row index set of active constraints of real-time clearing
at time $\tau$ on day $d$.
Variables
$\bm{p}_{d,\tau}$ Traditional generator day-ahead schedule at time $\tau$ on day $d$. $\bm{w}_{d,\tau}$ Renewable generator day-ahead schedule at time $\tau$ on day $d$.
$\bm{p}_{d,\tau}^{+}$
Traditional generator real-time adjustment for up-regulation
at time $\tau$ on day $d$.
$\bm{p}_{d,\tau}^{-}$
Traditional generator real-time adjustment for down-regulation
at time $\tau$ on day $d$.
$\bm{\kappa}_{d,\tau}$ Renewable generation spill at time $\tau$ on day $d$. $\bm{x}_{d}$ Collection of day-ahead decisions on day $d$.
$\bm{z}_{d,\tau}$ Collection of real-time decisions at time $\tau$ on day $d$. $\Theta$ Forecasting model parameter.
$\bm{p}_{d,\tau}^{+-}$
Collection of real-time up-/down- regulation adjustment
at time $\tau$ on day $d$.
$\bm{\sigma}_{d}$ day-ahead dual decisions on day $d$.
$\bm{\nu}_{d,\tau}$
Real-time dual decision at time $\tau=1$ on day $d$.
$\bm{\xi}_{d,\tau}$
Real-time dual decision at time $\tau=2,...,T$ on day $d$.
Parameters
$\bm{\rho}$ Marginal day-ahead generation cost. $\bm{l}_{d,\tau}$ Electricity load at time $\tau$ on day $d$.
$\overline{\bm{f}}$ Transmission line capacity $\overline{\bm{p}}$ Max day-ahead output of traditional generator
$\overline{\bm{r}}$ Max ramping capacity of traditional generator $\bm{\rho}_{+}$ Marginal real-time generation cost for up-regulation.
$\bm{\rho}_{-}$ Marginal real-time generation utility for down-regulation. $\bm{y}_{d,\tau}$ Renewable generation realization
$\bm{H}$ Power transfer distribution factors $\bm{s}_{d,\tau}$ Forecasting context
$\overline{\bm{y}}_{d,\tau}$ Renewable energy capacity
@@ II-A Sequential Market Clearing
We consider the sequential clearing of DA and RT markets [1]. The DA market is cleared at time $t$ on day d1d-1, with an advance of $k$ hours in time to the next day $d$, and covers energy transactions on day $d$, typically on an hourly basis. RES production is uncertain during the DA market, where the RES forecast acts as the RIEQ. Due to inevitable forecasting errors, the energy imbalance caused by forecasting errors needs to be settled in the RT market. Concretely, in the DA market, the operator determines the schedules of generators and RES to satisfy inelastic demand. The generation and RES schedule for each time-slot $\tau,\forall\tau=1,...,24$ on the next day $d$ are denoted as $\bm{p}_{d,\tau}$ and $\bm{w}_{d,\tau}$, respectively. The DA market clearing is,
@@
s.t. $\bm{1}^{\top}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau})=\bm{1}^{\top}\bm{l}_{d,\tau}:\gamma_{d,\tau},\forall\tau=1,...,T$ (1b)
$-\overline{\bm{f}}\leq\bm{H}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau}-\bm{l}_{d,\tau})\leq\overline{\bm{f}}:\underline{\mu}_{d,\tau},\overline{\mu}_{d,\tau},$
$\forall\tau=1,...,T$ (1c)
$0\leq\bm{p}_{d,\tau}\leq\overline{\bm{p}},\forall\tau=1,...,T$ (1d)
$-\overline{\bm{r}}\leq\bm{p}_{d,\tau}-\bm{p}_{d,\tau-1}\leq\overline{\bm{r}},\forall\tau=2,...,T$ (1e)
@@
In line with European practices [12, 20], we do not incorporate binary decisions regarding unit commitment (UC) in (1). However, we note that UC is a requisite consideration in the U.S. markets. To analyze the market behavior, the relaxed UC problem, where the binary commitment decisions are substituted with continuous ones, is widely used [3, 8]. The exact relaxation can be achieved by accurately adding cutting planes [21]. In this way, the mixed integer program is equivalently transformed into a linear program. A common practice is to train the forecasting model using the relaxed UC formulation [8]. During the operational phase, the generated forecasts are then used as inputs to the UC problem.
Since the RES production is uncertain in DA, the DA schedules are to be adjusted at each time-slot $\tau,\forall\tau=1,...,T$ in RT on day $d$, after the RES realization $\bm{y}_{d,\tau}$ is observed. The RT market deals with the imbalance $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ caused by RES, with a minimized imbalance cost. Additionally, the RT market clearing at time-slot $\tau,\forall\tau=2,...,T$ is influenced not only by the DA clearing outcomes but also by the extent of power adjustments made in the preceding time-slot. This is due to the ramping constraints that interconnect adjacent time slots. Because the market clearing is conducted separately for each day, the RT market clearing at time-slot $\tau=1$ on day $d$ remains unaffected by any adjustments made at time-slot $\tau=T$ on the previous day d1d-1. In the following, we firstly give the mathematical formulation of the RT clearing at $\tau=1$,
@@
In this work, we follow the practice in [7] and assume that the DA and RT markets share the same temporal granularity, i.e., one hour, leading to $T=24$. However, the market clearing model in (1), (2), and (3) can be easily adapted to DA and RT markets with different temporal granularity.
@@ II-B Mathematical Reformulation
$-\bm{p}_{d,\tau}^{*}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{p}}-\bm{p}_{d,\tau}^{*}$ (5a)
$\bm{p}_{d,\tau}^{*}-\overline{\bm{p}}\leq\bm{p}_{d,\tau}^{-}\leq\bm{p}_{d,\tau}^{*}$ (5b)
Since $0\leq\bm{p}_{d,\tau}^{*}\leq\overline{\bm{p}}$, the left side of (5) is less than 0. Considering the power adjustment $\bm{p}_{d,\tau}^{+},\bm{p}_{d,\tau}^{-}$ is larger than 0 as stated in (2d) and (2e), (5) can be further simplified as,
@@
The market clearing model developed in this work accounts for the participation of flexible conventional generators and RES on the supply side. This modeling approach is consistent with those used in [7, 8]. Although the discussions in the following sections are based on the DA and RT models given in (1), (2), and (3), we emphasize that even when additional participants such as energy storage are included, provided the model remains linear and the structure of the DA and RT market clearing problems still aligns with the formulations in (10), (11), and (12), the proposed method for deriving the loss function remains applicable. In the Appendix A, we use energy storage as an example to show that the structure of the market clearing model with energy storage aligns with (10), (11) (12).
III Methodology
Figure 1: An illustration of the parameter estimation problem.
[…]
@@ III Methodology
where $\hat{\bm{y}}_{d,\tau}$ denotes the forecast.
Typically, there are two phases, i.e., the training phase and the operational phase. During the training phase, the parameter $\Theta$ is estimated using historical data by solving an optimization problem in which the loss function serves as the objective. When a neural network is used as the forecasting model, this estimation involves forward computation followed by backward gradient descent. At the operational phase, forecasts are generated using the estimated parameters through a computationally efficient forward pass..
In the following, we first formulate the training phase as a bilevel optimization problem [16] to estimate the parameters of the forecasting model, followed by the operational phase. We then discuss key market properties under the issued RIEQ.
III-A Training Phase
Data in the training set $\{\{\bm{s}_{d,\tau},y_{d,\tau}\}_{\tau=1}^{T}\}_{d=1}^{D}$ is available, which consists of historical context and RES realization in $D$ days. An illustration of the bilevel program is shown in Fig. 1. The objective of determining RIEQ is to minimize the overall DA and RT operating costs while accounting for the impact of DA market decisions made under uncertainty on RT market clearing, which corrects power imbalances caused by RES. As a result, even with a deterministic DA clearing, the impact of RES uncertainty on both DA and RT markets is accounted for through RIEQ, ensuring coordination between the two markets. The upper level determines the model parameter $\Theta$, while the lower level involves the DA and RT market clearings. The bilevel program is mathematically formulated as,
@@ III-A Training Phase
[…]
where (15d) is the dual problem of DA clearing. (15f) is the dual problem of RT clearing at $\tau=1$, and (15h) is the dual problem of RT clearing at $\tau=2,...,T$. Since RT clearing requires the primal solutions of DA clearing and the previous RT clearing as input parameters, we also include the primal problems in the lower level. The forecast $\hat{\bm{y}}_{d}$ affects the upper-level objective (15a) via its impact on the DA and RT dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*}$, and their primal solutions $\bm{x}^{*}_{d,\tau},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$. If we can obtain the function between them and the forecast $\hat{\bm{y}}_{d}$ directly, the upper-level objective can be rewritten as a function regarding the forecast $\hat{\bm{y}}_{d}$, and can be used as the loss function for training. The specific design of the loss function is deferred to Section IV.
We note that the training is conducted offline. That is, it is performed using historical data from DA and RT markets, including demand, RES realizations, and generator bidding information. No actual market clearing occurs during the training phase.
III-B Operational Phase
At the operational phase, the forecast $\bm{\hat{y}}_{d,\tau}$ under the context $\bm{s}_{d,\tau}$ can be obtained by the trained model. Subsequently, utilizing this forecast, the operator proceeds to solve the DA clearing in (1), and obtains the DA schedules of traditional generator and RES, i.e., $\bm{p}_{d,\tau}^{*},\bm{w}_{d,\tau}^{*},\forall\tau=1,...,T$. After the RES realization is revealed at each time $\tau=1,\ldots,T$, the RT market described in (2) and (3) is cleared to resolve the power imbalance between the realized RES output and its scheduled value, i.e., $\bm{1}^{\top}(\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*})$.
III-C Market Properties
In this work, we assume that the DA market is primarily affected by uncertainty, while the RT market is cleared based on RES realization or a highly accurate RT RES forecast close to the realization. To improve the coordination between DA and RT markets, RIEQ appears as the DA RES forecast $\hat{\bm{y}}_{d,\tau}$ in (1f). Compared to the existing DA market clearing model that uses traditional DA RES forecasts, such as the expected RES generation, the market clearing structure remains unchanged and the values of RES forecasts is the only difference. In this regard, the current DA market properties are preserved, i.e.,
1) Cost recovery: The profits of market players are nonnegative at the market clearing solutions. As the DA market clearing in (1) determines the schedules of traditional generators and RES, cost recovery specifically ensures that their profits remain nonnegative given the market clearing schedules.
2) Revenue adequacy: Revenue adequacy implies that, at the market clearing solutions, the total payment made by the load to the market operator equals the total payment made by the market operator to traditional and renewable generators, as well as to the transmission line operator.
The proof for the above properties is provided in the Appendix B. Note that these properties can only be ensured in expectation for stochastic market clearing, which is a key drawback.